# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

### v1

@@ III Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{\alpha^{2}\left(M_{K}^{2}-3m_{s}^{*2}\right)}{2M_{K}\left(m_{s}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{3\left(m_{s}^{*}+M_{\eta}\right)}-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{s}^{*}+M_{\eta^{\prime}}\right)}\right]\mu_{s}^{*}.$
@@
In the present study, we have included the effect of relativistic quark confinement and have calculated the mass adjusted effective magnetic moments of its constituent quarks by the expression fiftyfive
$\mu_{d}^{*}=-\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{s}^{*}=-\frac{m_{u}^{*}}{m_{s}^{*}}\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{u}^{*}=-2\mu_{d}^{*}.$ (46)
In the above relations, $\Delta M=M_{B,vac}-M_{B}^{*}$ where $M_{B,vac}$ is the vacuum mass of the given baryon and $M_{B}^{*}$ denotes the in-medium mass of this baryon.
In light of this formalism, we can obtain the spin contribution ($\mu_{B}^{*S}$) towards the baryonic magnetic moment ($\mu^{*}_{B}$). The expressions for spin contribution of n* resonance, of $\frac{1}{2}^{-}$ baryon octet, is obtained as in thirtyone
$\mu_{\text{val}}^{*S}\left(n^{*}\right)=-\left(\frac{2}{9}\mu_{d}^{*}+\frac{1}{9}\mu_{u}^{*}\right)\cos^{2}\theta+\left(\frac{10}{9}\mu_{d}^{*}+\frac{5}{9}\mu_{u}^{*}\right)\sin^{2}\theta\\ -\frac{8}{9}\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
$\hskip 15.6491pt\mu_{\text{sea }}^{*S}\left(n^{*}\right)=$ $\frac{a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\mu_{d}^{*}\right.\left.+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\cos^{2}\theta$
$-\frac{5a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\right.\left.\mu_{d}^{*}+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\sin^{2}\theta$
$+\frac{8a}{9}\left(1+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
$\mu_{\text{orbit }}^{*S}\left(n^{*}\right)=$ $-\left(\frac{2}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{1}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\cos^{2}\theta+\left(\frac{10}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{5}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin^{2}\theta$ (47)
$-\frac{8}{9}\left(\mu^{*}\left(d_{\uparrow}\rightarrow\right)-\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin\theta\cos\theta.$
To account for the contributions of orbital angular momentum polarizations ($\mu^{*L}_{B}$) towards the total baryonic magnetic moment, we have calculated the individual contributions from the valence quarks and quark sea as
@@
$q^{(\pm 1)}\rightarrow-T_{[q,GB]}q^{(\pm 1)}+\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}.\hskip 150.79968pt$ (49)
@@
In eq. 49, $T_{[q,GB]}$ refers to the emission probability of a goldstone bosons (GB) from a quark $q^{(\pm 1)}$ and $\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}$ represents the transformation probability of a $q^{(\pm 1)}$ quark into two other quarks of $SU(3)$ model, given in thirtyone as
$\left|\psi\left(u^{(\pm 1)}\right)\right|^{2}=ad^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$ (50)
$\left|\psi\left(d^{(\pm 1)}\right)\right|^{2}=au^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$
$\left|\psi\left(s^{(\pm 1)}\right)\right|^{2}=a\alpha^{2}u^{(\pm 1)}+a\alpha^{2}d^{(\pm 1)}.$
The probability of transition a $q^{\pm 1}$ constituent quark into all the possible quarks ($u,d$ and $s$) along with GBs $\left(q\bar{q}^{\prime}\right)$ can be expressed as fiftyeight
$P\left(u^{\pm}\rightarrow u\bar{q}^{\prime}+q^{\prime\mp}\right)=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\mp}+ad^{\mp}+a\alpha^{2}s^{\mp},$ (51)
$P\left(d^{\pm}\rightarrow d\bar{q}^{\prime}+q^{\prime\mp}\right)=au^{\mp}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\mp}+a\alpha^{2}s^{\mp},$
$P\left(s^{\pm}\rightarrow s\bar{q}^{\prime}+q^{\prime\mp}\right)=a\alpha^{2}u^{\mp}+a\alpha^{2}d^{\mp}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\mp}.$
Using eq. 51, we can easily obtain the $T_{[q,GB]}$ for each constituent quark. The valence and sea quark contributions of the orbital angular momentum part of the magnetic moment of the baryons, can thus be calculated easily. As an example, we have presented the orbital angular momentum contribution to baryonic magnetic moment of $\frac{1}{2}^{-}$ $n^{*}$ resonances as,
$\mu^{*L}_{\mathrm{valence}}\left(n^{*}\right)=$ $\left(\frac{4}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\cos^{2}\theta-\left(\frac{1}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\sin^{2}\theta,$ (52)
$\mu^{*L}_{\mathrm{sea}}\left(n^{*}\right)=$ $\frac{2a}{9}\left[-\left(1+2\alpha^{2}\right)\mu^{*}_{d}+\left(1-\alpha^{2}\right)\mu^{*}_{u}\right.\left.+3\alpha^{2}\mu^{*}_{s}\right]\cos^{2}\theta$
$+\frac{a}{9}\left[-\left(1-\alpha^{2}\right)\mu^{*}_{d}\right.\left.+\left(1+2\alpha^{2}\right)\mu^{*}_{u}-3\alpha^{2}\mu^{*}_{s}\right]\sin^{2}\theta.$
Using similar methodology as demonstrated above, the magnetic moments of the $\frac{1}{2}^{-}$ octet baryon resonances $p^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$ and $\Xi^{*0}$ as well as the low-lying $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$, and $S_{11}^{0}(1650)$ have been calculated within the $\chi\mathrm{CQM}$ framework.
IV Results & Discussion
@@
We now discuss numerical results on the in-medium masses (section IV.1) and magnetic moments (section IV.2) of spin $\frac{1}{2}^{-}$ octet baryon resonances. The effective baryon masses have been listed in tables 1 and 2, so as to facilitate a discussion on the mass modifications. As an outcome of the calculations discussed in the previous sections, the vacuum and in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances have been illustrated in tables 3, 4, 5 and 6. For the sake of comparison with the baryonic magnetic moments obtained from earlier studies, we have tabulated the values calculated through different models in table 7 for $\rho_{B}=0$.
IV.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
Figure 1: Effective masses of constituent quarks $u,d$ and $s$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
As the medium modifications of baryon masses is calculated using the medium modified masses of constituent quarks, we first discuss the impact of density and temperature on effective masses of quarks. In fig. 1, the in-medium masses, $m^{*}_{u},m^{*}_{d}$ and $m^{*}_{s}$, of the constituent quarks are shown as a function of density in units of nuclear saturation density, at $T=0$ MeV (on the left) and $T=100$ MeV (on the right)(where nuclear saturation density $\rho_{0}=0.16$ $\text{fm}^{-3}$). It also illustrates the in-medium modification in quark masses for different values of isospin asymmetry parameter ($i_{a}$) and strangeness fraction ($f_{s}$).
@@ IV.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
From fig. 1 it is observed that as we move from vacuum ($\rho_{B}=0$) to denser hadronic medium, the quark masses decrease monotonously. For symmetric nuclear matter at temperature $T=0$ MeV, the effective masses of light quarks decreases by 71.14% and that of $s$ quark decreases by 14.65%, at density value of $\rho_{B}=3\rho_{0}$. As observed, the effects of density on the $s$ quark is significantly lower than that on the light quarks, which is an implication to the negligible coupling of $s$ quark with $\sigma$ field fiftyfour ; sixtytwo . By introducing finite strangeness in the symmetric nuclear medium, at $f_{s}=0.3$ and $T=0$ MeV, with an increase in baryon density there is an appreciable drop in $m^{*}_{s}$, relative to that in symmetric nuclear matter, while in the light quarks only slight variations in the $m^{*}_{u}$ and $m^{*}_{d}$ values are observed. It is also observed that with the rise in temperature at $T=100$ MeV the effective masses of $u$ and $d$ quarks rises by 8.91 % at $\rho_{B}=\rho_{0}$ and the $s$ quark mass elevates by just 1.38 %. Thus an increase in the medium temperature causes a rise in the masses of constituent quarks, at the same density. The increase in isospin asymmetry of the medium leads to increase in-medium masses of constituent quarks. A study on the basis of an extended NJL model has previously demonstrated that while the in-medium masses of light quarks approach their current quark masses, the $s$ quark condensate remains large even at densities close to $\rho_{B}\approx 1$ $\text{fm}^{-3}$ xia .
@@
In figs. 2, 3 and 4 and fig. 5 we have shown the variations in effective baryonic masses, as a function of medium density, for $\frac{1}{2}^{-}$ octet baryon resonances and $N^{*}$ resonance $S^{+(0)}_{11}(1650)$, respectively. Using eq. 25 the effective masses of $\frac{1}{2}^{-}$ baryon resonances are calculated in the isospin asymmetric strange matter, where the $E_{ispin}$ values are best fitted to procure the masses of $\frac{1}{2}^{-}$ baryon resonances in free space. A quick glance at the arguments and figures depicts, a monotonous decrease in the in-medium baryon masses as a function of density. The baryon masses in the medium fall rapidly in the lower density region up to $\rho_{B}=\rho_{0}$ and then saturate as the medium approaches large densities. As seen from fig. 2, when symmetric nuclear matter density rises upto $\rho_{B}=3\rho_{0}$, the mass of $p^{*}$ resonance drops appreciably by 33.2%. At a density of $3\rho_{0}$, with finite isospin asymmetry $i_{a}=0.3$, the mass of $p^{*}$ is 1010.03 MeV which is a rise from 1003.60 MeV at $i_{a}=0$. In fact, with higher valence $d$ quark content in a baryon, the density effects are found to be suppressed. This can be seen from the 25.27% decrease in the effective mass of $n^{*}$, as the medium density reaches $\rho_{B}=3\rho_{0}$ from vacuum. Keeping the density unchanged, say at $\rho_{B}=3\rho_{0}$, the $M^{*}_{n}$ value is found to incur a marginal increase, as a function of asymmetry of the medium, to attain 1149.18(1147.02) MeV for $i_{a}=0.3(0)$. In contrary to light quark content baryons, for symmetric nuclear matter the mass modifications incurred by $\Sigma^{*}$ resonances, as seen in fig. 3, is 18.5% and that of $\Xi^{*}$ resonances in fig. 4 is 14.5%. We can clearly state that, the density dependency on the baryons constituted by light quarks is much higher than that of hyperons, which is because $m^{*}_{u}$ and $m^{*}_{d}$ are more sensitive to density changes as compared to $m^{*}_{s}$. Additionally in hyperons, the in-medium effects are subject to the extent of $s$ quarks constituting a baryon. This is because as the number of constituent $s$ quarks increase, the in-medium effects on the constituent quark masses start having a dominating influence on the total mass of the baryon sixtytwo . In fig. 5, the masses of $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ varying as a function of medium density, have been shown. The extent of in-medium variations of the $N^{*}$ resonances is large compared to hyperons, like, in the case of $p^{*}$ and $n^{*}$ resonances. At a given value of $\rho_{B}$, with $i_{a}=0,T=0$ MeV, the impact of strangeness fraction, $f_{s}$, causes a sizable decrease in the masses of baryons. Looking up the mass of $\Xi^{*-}$ from table 1, at $\rho_{B}=3\rho_{0},f_{s}=0.3(0)$, we observe that the value decreases to 1399.14(1439.75) MeV. A cursory look at the plots of effective baryonic masses versus medium density suggests that, if the temperature of the medium increases to $T=100$ MeV then the mass modification curves will have relatively less steeper slope than that at $T=0$ MeV. For example, at $i_{a}=0,f_{s}=0(0.3)$ when the medium density is at $\rho_{B}=3\rho_{0}$, the effective mass of $\Sigma^{*+}$ being of 1300.88(1276.30) MeV, in table 1, at $T=0$ MeV increases to 1322.58(1287.69) MeV, in table 2, at $T=100$ MeV.
@@
Figure 2: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
@@
Figure 3: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 4: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 5: Effective masses of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
IV.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances

### v2

@@ 2.2 Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{\alpha^{2}}{2}\frac{\left(M_{K}^{2}-3m_{s}^{*2}\right)}{M_{K}\left(m_{s}^{*}+M_{K}\right)}-\frac{\beta^{2}}{3}\frac{M_{\eta}}{\left(m_{s}^{*}+M_{\eta}\right)}-\frac{\zeta^{\prime 2}}{6}\frac{M_{\eta^{\prime}}}{\left(m_{s}^{*}+M_{\eta^{\prime}}\right)}\right]\mu_{s}^{*}.$
@@
In the present study, we have included the effect of relativistic quark confinement and have calculated the mass adjusted in-medium magnetic moments of its constituent quarks by the expression fiftyfive
$\mu_{u}^{*}=2\left(1-\frac{\Delta M_{B}}{M_{B}^{*}}\right),\quad\mu_{d}^{*}=-\frac{1}{2}\mu_{u}^{*},\quad\mu_{s}^{*}=-\frac{m_{u}^{*}}{m_{s}^{*}}\left(1-\frac{\Delta M_{B}}{M_{B}^{*}}\right).$ (40)
In the above relations, $\Delta M_{B}=M_{B,vac}-M_{B}^{*}$. Here, $M_{B,vac}$ and $M_{B}^{*}$ denotes the vacuum and in-medium mass of a given baryon, respectively. As discussed in the introduction of present manuscript, in Ref. thirtyone the impact of relativistic quark confinement was not considered and the simplified expression $\mu_{q}=\frac{e_{q}}{2m_{q}},q=u,d,s$ was used to calculate the magnetic moment of constituent quarks in the free space. To account for the contributions of orbital angular momentum polarizations ($\mu^{*L}_{B}$) towards the total baryonic magnetic moment, we have calculated the individual contributions from the valence quarks and quark sea as
@@
$q^{(\pm 1)}\rightarrow-T_{[q,NGB]}q^{(\pm 1)}+\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}.\hskip 150.79968pt$ (42)
@@
In eq. 42, $T_{[q,NGB]}$ refers to the emission probability of a Nambu-Goldstone bosons (NGB) from a quark $q^{(\pm 1)}$ and $\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}$ represents the transformation probability of a $q^{(\pm 1)}$ quark into two other quarks of $SU(3)$ model, as given in Ref. thirtyone .
Using the probability of transition of a constituent quark ($q^{\pm 1}$) into all other quarks ($u,d$ and $s$) accompanied by NGBs fiftyeight , we can easily obtain the $T_{[q,NGB]}$ for each constituent quark. The contributed values from the valence and sea quark polarizations of the orbital angular momentum part of the magnetic moment of the baryons, can thus be calculated easily. Working on the framework as described in this section, the magnetic moments of the $\frac{1}{2}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$ and $\Xi^{*0}$ as well as the low-lying $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$, and $S_{11}^{0}(1650)$ have been calculated within the $\chi\mathrm{CQM}$ framework.
3 Results & Discussion
@@
We now discuss numerical results on the in-medium masses (section 3.1) and magnetic moments (section 3.2) of spin $\frac{1}{2}^{-}$ octet baryon resonances. Various parameters used in the current manuscript are listed in Table 1. The effective baryon masses have been listed in Tables 2 and 3, so as to facilitate a discussion on the mass modifications. As an outcome of the calculations discussed in the previous sections, the vacuum and in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances have been illustrated in Tables 4 to 7. For the sake of comparison with the baryonic magnetic moments obtained from earlier studies, we have tabulated the values calculated through different models in Table 8 for $\rho_{B}=0$. As discussed in previous section, in our present work quark masses and energies are defined in terms of scalar and vector fields. The values of the confining parameter $k_{c}$ and couplings are fitted to binding energy value of $-16$ MeV at the nuclear saturation density. With this approach, at $\rho_{B}=0$, the constituent quark masses of $u$ and $d$ quarks are obtained to be 256 MeV while for s-quark it is 477 MeV. The difference in the constituent quark masses in present manuscript compared to Ref. thirtyone (for $u/d$ quarks 330 MeV and for $s$ quark 490 MeV), leads to different values for the magnetic moments of baryons in the free space as can be seen in Table 8.
3.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
Figure 1: In-medium masses of constituent quarks $u,d$ and $s$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
@@ 3.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
As the in-medium baryon masses is determined using the medium modified masses of constituent quarks, we first discuss the impact of density and temperature on the effective quark masses. In fig. 1, the constituent quark masses, $m^{*}_{u},m^{*}_{d}$ and $m^{*}_{s}$, are shown to vary with the relative hadronic matter density $\rho_{B}/\rho_{0}$, at $T=0$ MeV (on the left) and $T=100$ MeV (on the right)(where nuclear saturation density $\rho_{0}=0.16$ $\text{fm}^{-3}$). It also illustrates the in-medium mass modifications of quarks, subject to a medium with finite isospin asymmetry ($i_{a}$) and strangeness fraction ($f_{s}$). From fig. 1 it is observed that as we move from vacuum ($\rho_{B}=0$) to denser hadronic medium, the quark masses decrease monotonously. For symmetric nuclear matter at temperature $T=0$ MeV, the effective masses of light quarks decreases by 71.14% and that of $s$ quark decreases by 14.65%, at density value of $\rho_{B}=3\rho_{0}$. As observed, the effects of density on the $s$ quark is significantly lower than that on the light quarks, which is an implication to the negligible coupling of $s$ quark with $\sigma$ field fiftyfour ; sixtytwo . By introducing finite strangeness in the symmetric nuclear medium, at $f_{s}=0.3$ and $T=0$ MeV, with an increase in baryon density there is an appreciable drop in $m^{*}_{s}$, relative to that in symmetric nuclear matter, while in the light quarks only slight variations in the $m^{*}_{u}$ and $m^{*}_{d}$ values are observed. It is also observed that with the rise in temperature at $T=100$ MeV the effective masses of $u$ and $d$ quarks rises by 8.91 % at $\rho_{B}=\rho_{0}$ and the $s$ quark mass elevates by just 1.38 %. Thus an increase in the medium temperature causes a rise in the masses of constituent quarks, at the same density. Keeping the medium density affixed, a rise in isospin asymmetry parameter leads to a surge in-medium masses of constituent quarks. A study on the basis of an extended NJL model has previously demonstrated that while the in-medium masses of light quarks approach their current quark masses, the $s$ quark condensate remains large even at densities close to $\rho_{B}\approx 1$ $\text{fm}^{-3}$ xia .
@@
In figs. 2, 3 and 4 and fig. 5 we have shown the variations in effective baryonic masses, as a function of medium density, for $\frac{1}{2}^{-}$ octet baryon resonances and $N^{*}$ resonance $S^{+(0)}_{11}(1650)$, respectively. Using eq. 26 the in-medium masses of $\frac{1}{2}^{-}$ baryon resonances are computed in the presence of a hadronic matter, where the $E_{ispin}$ values are best fitted to procure the masses of $\frac{1}{2}^{-}$ baryon resonances in free space. A quick glance at the arguments and figures depicts, a monotonous decrease in the in-medium baryon masses as a function of density. The baryon masses in the medium fall rapidly in the lower density region up to $\rho_{B}=\rho_{0}$ and then saturate as the medium approaches large densities. As seen from fig. 2, when symmetric nuclear matter density rises upto $\rho_{B}=3\rho_{0}$, the mass of $p^{*}$ resonance drops appreciably by a decrease of 25.27%. At a density of $3\rho_{0}$, with finite isospin asymmetry $i_{a}=0.3$, the mass of $p^{*}$ is 1149.18 MeV which is a mild rise from 1147.02 MeV at $i_{a}=0$. Taking into account a temperature rise from $T=0$ MeV to 100 MeV, we see distinctive baryonic mass curves for different strangeness fractions ($f_{s}=0,0.3$) of the medium. This exact same behaviour is seen to be replicated by the medium modified mass plots of $n^{*}$ as can be seen from fig. 2, which is quite trivial as well considering the light quark compositions of $p^{*}$ and $n^{*}$. In contrary to light quark content baryons, for symmetric nuclear matter the mass modifications incurred by $\Sigma^{*}$ resonances, as seen in fig. 3, is 18.5% and that of $\Xi^{*}$ resonances in fig. 4 is 14.5%. We can clearly state that, the density dependency on the baryons constituted by light quarks is much higher than that of hyperons, which is because $m^{*}_{u}$ and $m^{*}_{d}$ are more sensitive to density changes as compared to $m^{*}_{s}$. Additionally in hyperons, the in-medium effects are subject to the extent of $s$ quarks constituting a baryon. This is because as the number of constituent $s$ quarks increase, the in-medium effects on the constituent quark masses start having a dominating influence on the total mass of the baryon sixtytwo . In fig. 5, the masses of $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ varying as a function of medium density, have been shown. The extent of in-medium variations of the $N^{*}$ resonances is large compared to hyperons, like, in the case of $p^{*}$ and $n^{*}$ resonances. At a given value of $\rho_{B}$, with $i_{a}=0,T=0$ MeV, the influence of non-zero strangeness fraction, $f_{s}$, leads to a sizable decrease in the masses of baryons. Looking up the mass of $\Xi^{*-}$ from table 2, at $\rho_{B}=3\rho_{0},f_{s}=0.3(0)$, we observe that the value decreases to 1399.14(1439.75) MeV. A cursory look at the plots of effective baryonic masses versus medium density suggests that, if the temperature of the medium increases to $T=100$ MeV then the mass modification curves will have relatively less steeper slope than that at $T=0$ MeV. For example, at $i_{a}=0,f_{s}=0(0.3)$ when the medium density is at $\rho_{B}=3\rho_{0}$, the effective mass of $\Sigma^{*+}$ being of 1300.88(1276.30) MeV, in table 2, at $T=0$ MeV increases to 1322.58(1287.69) MeV, in table 3, at $T=100$ MeV.
@@
Figure 2: In-medium masses of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
@@
Figure 3: In-medium masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
Figure 4: In-medium masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
Figure 5: In-medium masses of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
3.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances